# envname-P envname#1

### v2

@@ 4.1 Experimental setup
Following prior work Shi et al. (2025), we evaluate both the compressor and the end-to-end framework on the full development sets of four widely used question answering benchmarks: NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultihopQA (Ho et al., 2020).
Models CR NQ HotpotQA Musique 2Wiki Average
@@
Autocompressor 1x 17.24 14.61 3.81 19.89 13.89
Mistral-7B w/o BGE retrieval 1x 35.01 27.55 5.38 38.45 26.6
Mistral-7B w/ BGE retrieval 1x 54.58 42.94 8.94 44.24 37.67
Phi4-mini w/o BGE retrieval 1x 18.77 21.10 4.05 30.26 18.55
Phi4-mini w/ BGE retrieval 1x 48.14 37.78 8.11 35.11 32.28
llmlingua-2 4x 47.53 37.05 9.02 44.35 34.49
SCP-Mistral-7B 4x 57.05+9.52 45.09+8.04 10.34+1.32 46.94+2.59 39.86+5.37
SCP-Phi4-mini 4x 53.31+5.78 42.36+5.31 8.73-0.29 45.22+0.87 37.40+2.91
coconum 16x 24.12 21.48 3.52 24.48 18.40
pcc 16x 31.38 22.29 3.43 19.47 19.14
pisco 16x 54.39 41.94 10.09 44.88 37.83
SCP-Mistral-7B 16x 55.56+1.17 43.72+1.78 10.55+0.46 46.00+1.12 38.96+1.13
SCP-Phi4-mini 16x 51.96-2.43 40.86-1.08 8.61-1.48 44.27-0.61 36.42-1.42
xrag 128x 32.35 25.16 3.64 28.79 22.48
SCP-Mistral-7B 128x 53.36+21.01 41.37+16.21 10.26+6.62 46.40+17.61 37.85+15.37
SCP-Phi4-mini 128x 43.09+10.74 33.92+8.76 6.87+3.23 43.70+14.91 31.90+9.42
Oracle
Autocompressor 1x 29.47 19.24 7.16 26.74 20.65
Mistral-7B w/ BGE retrieval 1x 71.64 70.77 45.72 68.83 64.24
Phi4-mini w/ BGE retrieval 1x 66.10 64.06 37.07 52.69 54.98
llmlingua-2 4x 63.99 52.42 27.47 53.92 49.45
SCP-Mistral-7B 4x 76.50+12.51 73.81+21.39 46.26+18.79 70.48+16.56 66.76+17.31
SCP-Phi4-mini 4x 73.67+9.68 72.41+19.99 40.13+12.66 64.22+10.30 62.61+13.16
coconum 16x 25.61 21.72 3.64 24.63 18.90
pcc 16x 49.62 34.56 18.25 27.56 32.50
pisco 16x 73.44 66.53 33.80 60.45 58.55
SCP-Mistral-7B 16x 75.48+2.04 70.79+4.26 43.15+9.35 66.16+5.71 63.90+5.35
SCP-Phi4-mini 16x 73.17-0.27 70.26+3.73 38.39+4.59 63.15+2.70 61.24+2.69
xrag 128x 42.60 30.21 7.03 30.94 27.70
SCP-Mistral-7B 128x 69.96+27.36 62.09+31.88 30.86+23.83 59.08+28.14 55.50+27.80
SCP-Phi4-mini 128x 60.44+17.84 51.52+21.31 19.28+12.25 50.29+19.35 45.38+17.68
Table 2: Compressor performance on four QA datasets. The best performance is highlighted in bold. We show the absolute performance change (±) of our method under different compression rates relative to its corresponding best baseline performance. CR denotes compression rate.
@@
For compressor evaluation, we benchmark against both classical and recent methods, including AutoCompressor (Chevalier et al., 2023), XRAG(Cheng et al., 2025), COCOM (Rau et al., 2025), PCC (Dai et al., 2025), LLMLingual-2 (Pan et al., 2024), and PISCO (Louis et al., 2025a). For reranking, we compare with BM25, BGE-Reranker (Chen et al., 2023), RankZephyr-7B (Pradeep et al., 2023), Setwise (Zhuang et al., 2024), and Rank-R1 (Zhuang et al., 2025). End-to-end QA results are evaluated against representative RAG systems, including prompt-based (GenGround (Shi et al., 2024), In-Context RAG), retrieval-optimized (ReComp (Xu et al., 2024), DPA-RAG (Dong et al., 2025)), fine-tuned LLMs (Self-RAG (Asai et al., 2024), Retrobust (Yoran et al., 2024), ChatQA (Liu et al., 2025)), and jointly optimized models (DDR-RAG (Li et al., 2024a), DRO (Shi et al., 2025)). Unlike all baselines operating on raw text, our method is the first to jointly optimize reranking and generation directly over compressed representations. Full experimental settings are provided in Appendix B. Below, we summarize the key findings, while the complete set of additional experiments can be found in the Appendix, including pretraining data analysis (App.C & D), training process analysis (App.E), fidelity and grounding evaluations (App.F), as well as further module analyses (App.G).
@@ 4.2 Evaluation of Compression Effectiveness
We evaluate our document compressor under two settings: Normal and Oracle. In the Normal setting, the model retrieves the top-5 documents from Wikipedia-2021 for each query. In the Oracle setting, the annotated positive document is included among the top-5 to isolate compression quality from retrieval noise.
Table 2 summarizes results across compression ratios. For full results, please refer to table 6 Our method consistently outperforms all baselines. Compared to the best soft compression model PISCO, our model achieves average gains of 1.13% (Normal) and 5.35% (Oracle); over the hard compression baseline LLMLingual-2, improvements reach 5.37% and 17.31%, highlighting stronger semantic preservation.
Surprisingly, our model exceeds the text-based w/ BGE retrieval baseline using uncompressed documents, with average gains of 2.36% on Mistral-7B and 6.36% on Phi-4-min. This implies that well-trained soft compression can retain essential reasoning information while substantially reducing input length. This may be because the compressed representations filter out irrelevant content and focus the generator on the reasoning-relevant context, leading to better generalization than raw text inputs. While performance declines at extreme compression (beyond 32× in Oracle), the drop remains moderate under Normal conditions due to weaker document relevance.
Models CR NQ HotpotQA Musique 2Wiki Average
@@
Prompting-based Method
GenGround* 1x 42.31 40.60 44.71 41.27 24.36 20.77 42.58 39.61 38.49 35.56
In-context RAG* 1x 44.69 38.07 41.27 37.14 20.11 16.78 41.02 38.51 36.77 32.62
Retrieval tuning
RECOMP* 1x 42.67 37.47 42.72 38.72 24.96 17.34 38.26 32.17 37.15 31.43
DPA-RAG* 1x 44.31 37.29 40.53 37.15 20.36 18.45 39.66 39.02 36.22 32.98
LLM Fine-tuning
RetRobust* 1x 43.82 37.03 40.54 35.59 18.16 18.11 39.11 38.65 35.41 32.34
ChatQA* 1x 34.54 23.64 44.60 33.40 17.05 16.64 31.90 26.80 32.02 25.12
Self-RAG* 1x 31.63 29.74 27.30 16.30 21.50 9.43 27.33 23.52 26.94 19.75
End-to-end optimization
DDR-RAG* 1x 28.76 40.74 35.44 31.71 10.57 13.54 38.40 35.44 28.29 30.36
DRO-Mistral-7B* 1x 51.01 42.41 47.87 40.37 25.32 21.36 43.65 42.12 41.96 36.56
CLaRa-Mistral-7B (Pretrianing-initialized) 4x 40.62 31.21 39.53 29.54 14.53 6.16 42.59 38.49 34.32 26.35
16x 41.75 32.24 44.37 33.72 15.36 6.99 43.47 39.50 36.24 28.11
32x 40.68 31.36 41.84 31.26 15.32 6.66 43.23 38.98 35.27 27.06
CLaRa-Mistral-7B (Instruction-initialized) 4x 48.21 38.16 45.93 35.12 17.49 8.11 47.18 43.11 39.70 31.12
16x 50.89 41.02 47.62 36.67 18.01 8.44 44.66 40.48 40.30 31.65
32x 49.72 39.88 45.73 34.85 16.83 7.82 42.57 38.41 38.71 30.24
@@
CLaRa-Mistral-7B (Pretrianing-initialized) 4x 77.80 70.52 77.66 64.83 41.59 30.33 73.20 69.14 67.56 58.70
16x 73.81 65.74 69.57 56.76 31.15 21.18 65.90 61.31 60.11 51.25
32x 72.03 63.65 70.91 57.07 33.40 22.22 66.32 61.12 60.66 51.02
CLaRa-Mistral-7B (Instruction-initialized) 4x 75.63 67.64 69.66 56.92 33.19 22.42 73.86 69.74 63.08 54.18
16x 71.54 63.29 71.17 57.54 30.77 20.56 60.37 55.73 58.46 49.28
32x 69.75 65.17 68.87 55.20 28.87 18.45 64.38 59.32 57.97 49.53
Table 3: End-to-End QA Performance. * indicates results reported from the DRO paper. CR denotes compression rate. The highest scores are shown in bold, and the second-best ones are underlined. Overall, our method achieves comparable performance while reducing the required context length by 16×.
@@ 4.4 Retrieval performance
5 Ablation Study
Pretraining Data Mix
Each document in our setup is paired with two output types: (i) QA-style question–answer pairs and (ii) paraphrased documents. To assess the impact of data composition, we vary pretraining objectives and report results in Table 4 and 20. For both Mistral-7B and Phi4-mini, using either SimpleQA or Paraphrase alone already outperforms the no-pretraining baseline, showing that factual reasoning and paraphrastic rewriting both enrich compressed representations. Combining multiple QA types (SimpleQA+ComplexQA) or adding paraphrases (SimpleQA+ComplexQA+Para) achieves the best performance, confirming that diverse objectives enhance semantic coverage and generalization—especially under the Oracle setting, where high-quality retrieval amplifies pretraining benefits.
@@ 5 Ablation Study
Effect of MSE Loss
We analyze the effect of the mean-squared error (MSE) loss in Eq.2.2, which aligns compressed and original document representations. As shown in Table 5 and 21, the improvements are modest (0.3–0.6 points on average) but consistent across datasets, confirming that the MSE loss facilitates semantic preservation during compression. To provide a qualitative perspective, we visualize 4K document embeddings and their corresponding compressed representations using t-SNE (Figure 8 in Appendix). Without the MSE loss, the two distributions are clearly separated, reflecting a weak correspondence between the memory-token and document spaces. When the MSE loss is applied, the compressed embeddings exhibit strong overlap with the original document representations, demonstrating that the alignment objective effectively enforces semantic consistency between embedding spaces.
Models CR NQ HotpotQA Musique 2Wikiqa Average
Oracle
Mistral-7B 32x 74.65 69.05 37.32 62.98 61.00
   w/ mse 32x 73.77-0.88 69.51+0.46 38.31+0.99 64.54+1.56 61.53+0.53
Mistral-7B 128x 71.24 62.26 29.29 57.87 55.16
   w/ mse 128x 69.96-1.28 62.09-0.17 30.86+1.57 59.08+1.21 55.50+0.34
Table 5: Instruction-tuning performance with and without the MSE loss under different compression ratios (CR = 32, 128) and oracle retrieval settings.
@@ 6.2 End-to-End Optimization for Retrieval and Generation
As motivated earlier, joint training of retrieval and generation in RAG systems is hindered by the non-differentiability of discrete document selection. In typical QA pipelines, the retriever reorders retrieved documents before generation (Yu et al., 2024; Dong et al., 2024), but discrete sampling operations prevent gradient backpropagation. In contrast, our framework uniquely combines compression and joint training: by employing length-flexible compressed vectors in a shared latent space, we enable efficient differentiable selection while drastically reducing context length. Optimized solely through the generator’s language modeling loss, CLaRa ensures consistent training–inference alignment and efficient end-to-end learning without explicit retrieval supervision.
@@ 7 Conclusion
Limitations
Compressor Generalization.
The current compressor is pretrained exclusively on Wikipedia data. To improve generalizability, future work may incorporate domain-adaptive pretraining objectives and leverage more diverse corpora (e.g., code datasets (Wang et al., 2025)) to enhance representation robustness across modalities and domains.
Reasoning over Compressed Representations.
While our study does not primarily focuses on reasoning based on compressed representations, the compact and semantically dense nature of compressed representations makes them a promising candidate for integration into reasoning-oriented or agentic RAG frameworks, such as Search-R1 (Jin et al., 2025). A natural next step is to investigate whether such representations can function as efficient reasoning memory in multi-hop or planning-based RAG systems.
Model Size.
Our experiments are conducted on medium-scale models (Mistral-7B and Phi-4B). Larger models may produce higher-quality document representations that better capture semantic nuances. An open question for future work is whether there exists a model size threshold beyond which compressed representations surpass raw text in supporting understanding and generation.
Generalization of Implicit Representations.
This work focuses on leveraging unified representations for both understanding and generation within the RAG framework. Given the growing interest toward unified models that jointly perform comprehension and generation (Zhang et al., 2025), extending compression-based methods to broader tasks—such as tool learning (Qu et al., 2025)—offers a promising direction for developing more general-purpose, reasoning-capable systems. Besides, linking implicit understanding with implicit reasoning (Hao et al., 2025; Kang et al., 2025) would be an interesting direction.
@@
$\boxed{\;\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y\,|\,x)}\Big[\underbrace{p(d\,|\,x)\big(p(y\,|\,x,d)-p(y\,|\,x)\big)}_{\text{(I) probability path}}+\underbrace{p(d\,|\,x)\,\frac{\partial p(y\,|\,x,d)}{\partial s_{xd}}}_{\text{(II) representation/generation path}}\Big].\;}$ (A.7)
@@
$\boxed{\;\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y\,|\,x)}\;p(d\,|\,x)\,\big(p(y\,|\,x,d)-p(y\,|\,x)\big).\;}$ (A.9)
@@
$\boxed{\;\frac{\partial p(y\,|\,x,d)}{\partial s_{xd}}=p(y\,|\,x,r)\;g^{\top}\frac{\partial r}{\partial s_{xd}}=\frac{p(y\,|\,x,r)}{\tau}\;\pi_{d}\;g^{\top}\big(z_{d}-r\big).\;}$ (A.13)
@@
$\boxed{\;\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y\,|\,x)}\Big[p(d\,|\,x)\,\big(p(y\,|\,x,d)-p(y\,|\,x)\big)\;+\;\frac{p(r\,|\,x)\,p(y\,|\,x,r)}{\tau}\;\pi_{d}\;g^{\top}\big(z_{d}-r\big)\Big].\;}$ (A.14)

### v3

@@ 4.1 Experimental setup
We train the compressor on our curated synthetic data, perform compressor instruction tuning using the training data released by COCOM, and conduct end-to-end training on the corresponding training sets provided with the evaluation tasks. For evaluation, following prior work Shi et al. (2025), we evaluate both the compressor and the end-to-end framework on the full development sets of four widely used question answering benchmarks: NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultihopQA (Ho et al., 2020).
Table 1: Compressor performance on four QA datasets measured by CEM (%) (B.3). DF denotes document format. Compression rates (CR) is calculated as the ratio between the length of the raw text document and that of the compressed document. The best performance is highlighted in bold. We show the absolute performance change (±) of our method under different CR relative to its corresponding best baseline performance. For all methods where the LLM is not explicitly specified, Mistral-7B is used as the default generator.
Models CR DF W/ doc NQ HotpotQA Musique 2Wiki Average
@@
Autocompressor 1x Vectors $\checkmark$ 17.24 14.61 3.81 19.89 13.89
Mistral-7B 1x - $\times$ 35.01 27.55 5.38 38.45 26.6
Mistral-7B 1x Raw texts $\checkmark$ 54.58 42.94 8.94 44.24 37.67
Phi4-mini 1x - $\times$ 18.77 21.10 4.05 30.26 18.55
Phi4-mini 1x Raw texts $\checkmark$ 48.14 37.78 8.11 35.11 32.28
llmlingua-2 4x Raw texts $\checkmark$ 47.53 37.05 9.02 44.35 34.49
SCP-Mistral-7B 4x Vectors $\checkmark$ 57.05+9.52 45.09+8.04 10.34+1.32 46.94+2.59 39.86+5.37
SCP-Phi4-mini 4x Vectors $\checkmark$ 53.31+5.78 42.36+5.31 8.73-0.29 45.22+0.87 37.40+2.91
coconum 16x Vectors $\checkmark$ 24.12 21.48 3.52 24.48 18.40
pcc 16x Vectors $\checkmark$ 31.38 22.29 3.43 19.47 19.14
pisco 16x Vectors $\checkmark$ 54.39 41.94 10.09 44.88 37.83
SCP-Mistral-7B 16x Vectors $\checkmark$ 55.56+1.17 43.72+1.78 10.55+0.46 46.00+1.12 38.96+1.13
SCP-Phi4-mini 16x Vectors $\checkmark$ 51.96-2.43 40.86-1.08 8.61-1.48 44.27-0.61 36.42-1.42
xrag 128x Vectors $\checkmark$ 32.35 25.16 3.64 28.79 22.48
SCP-Mistral-7B 128x Vectors $\checkmark$ 53.36+21.01 41.37+16.21 10.26+6.62 46.40+17.61 37.85+15.37
SCP-Phi4-mini 128x Vectors $\checkmark$ 43.09+10.74 33.92+8.76 6.87+3.23 43.70+14.91 31.90+9.42
@@
For compressor evaluation, we benchmark against both classical and recent methods, including AutoCompressor (Chevalier et al., 2023), XRAG(Cheng et al., 2025), COCOM (Rau et al., 2025), PCC (Dai et al., 2025), LLMLingual-2 (Pan et al., 2024), and PISCO (Louis et al., 2025a). For reranking, we compare with BM25, BGE-Reranker (Chen et al., 2023), RankZephyr-7B (Pradeep et al., 2023), Setwise (Zhuang et al., 2024), and Rank-R1 (Zhuang et al., 2025). End-to-end QA results are evaluated against representative RAG systems, including prompt-based ( In-Context RAG), retrieval-optimized (ReComp (Xu et al., 2024), DPA-RAG (Dong et al., 2025)), fine-tuned LLMs (Self-RAG (Asai et al., 2024), Retrobust (Yoran et al., 2024), ChatQA (Liu et al., 2025), GenGround (Shi et al., 2024)), and jointly optimized models (DDR-RAG (Li et al., 2024a), DRO (Shi et al., 2025)). Unlike all baselines operating on raw text, our method is the first to jointly optimize reranking and generation directly over compressed representations. Full experimental settings are provided in App. B. Below, we summarize the key findings, while the complete set of additional experiments can be found in the Appendix, including pretraining data analysis (App. C & D), training process analysis (App. E), fidelity and grounding evaluations (App. F), as well as further module analyses (App. G).
@@ 4.2 Evaluation of Compression Effectiveness
We evaluate our document compressor under two settings: Normal and Oracle. In the Normal setting, for each test query, we use the BGE-large-en-v1.5 model to retrieve the top-5 documents from Wikipedia-2021. All baselines use the same set of retrieved context documents, which are then post-processed differently by each method. In the Oracle setting, the annotated positive document is included among the top-5 to isolate compression quality from retrieval noise. Table 1 summarizes results across compression ratios; see Table 5 for full results. Our method consistently outperforms all baselines. Compared to the best soft compression model PISCO, our model achieves average gains of 1.13% (Normal) and 5.35% (Oracle); over the hard compression baseline LLMLingual-2, improvements reach 5.37% and 17.31%, highlighting stronger semantic preservation.
Surprisingly, our model exceeds the text-based baseline using uncompressed documents, with average gains of 2.36% on Mistral-7B and 6.36% on Phi-4-min. This implies that well-trained soft compression can retain essential reasoning information while substantially reducing input length. This may be because the compressed representations filter out irrelevant content and focus the generator on the reasoning-relevant context, leading to better generalization than raw text inputs. While performance declines at extreme compression (beyond 32× in Oracle), the drop remains moderate under Normal conditions due to weaker document relevance.
Table 2: End-to-End QA Performance measured by EM (%) and F1 (%) (B.3). * indicates results reported from the DRO paper. CR denotes compression rate. The highest scores are shown in bold, and the second-best ones are underlined. Overall, our method achieves comparable performance while reducing the required context length by 16×.
Models CR Retriever tuning Generator tuning NQ HotpotQA Musique 2Wiki Average
@@
Normal data setting
In-context RAG* 1x $\times$ $\times$ 44.69 38.07 41.27 37.14 20.11 16.78 41.02 38.51 36.77 32.62
RECOMP* 1x $\checkmark$ $\times$ 42.67 37.47 42.72 38.72 24.96 17.34 38.26 32.17 37.15 31.43
DPA-RAG* 1x $\checkmark$ $\times$ 44.31 37.29 40.53 37.15 20.36 18.45 39.66 39.02 36.22 32.98
RetRobust* 1x $\times$ $\checkmark$ 43.82 37.03 40.54 35.59 18.16 18.11 39.11 38.65 35.41 32.34
ChatQA* 1x $\times$ $\checkmark$ 34.54 23.64 44.60 33.40 17.05 16.64 31.90 26.80 32.02 25.12
GenGround* 1x $\times$ $\checkmark$ 42.31 40.60 44.71 41.27 24.36 20.77 42.58 39.61 38.49 35.56
Self-RAG* 1x $\times$ $\checkmark$ 31.63 29.74 27.30 16.30 21.50 9.43 27.33 23.52 26.94 19.75
DDR-RAG* 1x $\checkmark$ $\checkmark$ 28.76 40.74 35.44 31.71 10.57 13.54 38.40 35.44 28.29 30.36
DRO-Mistral-7B* 1x $\checkmark$ $\checkmark$ 51.01 42.41 47.87 40.37 25.32 21.36 43.65 42.12 41.96 36.56
CLaRa-Mistral-7B (Pretraining-initialized) 4x $\checkmark$ $\checkmark$ 40.62 31.21 39.53 29.54 14.53 6.16 42.59 38.49 34.32 26.35
16x $\checkmark$ $\checkmark$ 41.75 32.24 44.37 33.72 15.36 6.99 43.47 39.50 36.24 28.11
32x $\checkmark$ $\checkmark$ 40.68 31.36 41.84 31.26 15.32 6.66 43.23 38.98 35.27 27.06
CLaRa-Mistral-7B (Instruction-initialized) 4x $\checkmark$ $\checkmark$ 48.21 38.16 45.93 35.12 17.49 8.11 47.18 43.11 39.70 31.12
16x $\checkmark$ $\checkmark$ 50.89 41.02 47.62 36.67 18.01 8.44 44.66 40.48 40.30 31.65
32x $\checkmark$ $\checkmark$ 49.72 39.88 45.73 34.85 16.83 7.82 42.57 38.41 38.71 30.24
@@
CLaRa-Mistral-7B (Pretraining-initialized) 4x $\checkmark$ $\checkmark$ 77.80 70.52 77.66 64.83 41.59 30.33 73.20 69.14 67.56 58.70
16x $\checkmark$ $\checkmark$ 73.81 65.74 69.57 56.76 31.15 21.18 65.90 61.31 60.11 51.25
32x $\checkmark$ $\checkmark$ 72.03 63.65 70.91 57.07 33.40 22.22 66.32 61.12 60.66 51.02
CLaRa-Mistral-7B (Instruction-initialized) 4x $\checkmark$ $\checkmark$ 75.63 67.64 69.66 56.92 33.19 22.42 73.86 69.74 63.08 54.18
16x $\checkmark$ $\checkmark$ 71.54 63.29 71.17 57.54 30.77 20.56 60.37 55.73 58.46 49.28
32x $\checkmark$ $\checkmark$ 69.75 65.17 68.87 55.20 28.87 18.45 64.38 59.32 57.97 49.53
@@ 4.4 Retrieval performance
5 Ablation Study
Pretraining Data Mix
Each document in our setup is paired with two output types: (i) QA-style question–answer pairs and (ii) paraphrased documents. To assess the impact of data composition, we vary pretraining objectives and report results in Tables 3 and 19. For both Mistral-7B and Phi4-mini, using either SimpleQA or Paraphrase alone already outperforms the no-pretraining baseline, showing that factual reasoning and paraphrastic rewriting both enrich compressed representations. Combining multiple QA types (SimpleQA+ComplexQA) or adding paraphrases (SimpleQA+ComplexQA+Para) achieves the best performance, confirming that diverse objectives enhance semantic coverage and generalization—especially under the Oracle setting, where high-quality retrieval amplifies pretraining benefits.
Effect of $\mathcal{L}_{\text{MSE}}$
Table 4: Instruction-tuning performance with and without MSE loss under different compression ratios and oracle retrieval settings.
Models CR NQ HotpotQA Musique 2Wikiqa
Mistral-7B 32x 74.65 69.05 37.32 62.98
   w/ mse 32x 73.77-0.88 69.51+0.46 38.31+0.99 64.54+1.56
Mistral-7B 128x 71.24 62.26 29.29 57.87
   w/ mse 128x 69.96-1.28 62.09-0.17 30.86+1.57 59.08+1.21
We analyze the effect of the MSE loss (Eq. 2.1), which aligns compressed and original document representations. As Tables 4 and 20 show, including this loss leads to a modest (0.3–0.6 points on average) but consistent improvement across datasets, confirming that it facilitates semantic preservation during compression. To provide a qualitative perspective, we visualize 4K document embeddings and their corresponding compressed representations using t-SNE (Fig 9 in Appendix). Without the MSE loss, the two distributions are clearly separated, reflecting a weak correspondence between the memory-token and document spaces. With the MSE loss, the compressed embeddings exhibit strong overlap with the original document representations, demonstrating that the alignment objective effectively enforces semantic consistency between embedding spaces.
@@ 6.2 End-to-End Optimization for RAG
As motivated earlier, joint training of retrieval and generation in RAG systems is hindered by the non-differentiability of discrete document selection. In typical QA pipelines, the retriever reorders retrieved documents before generation (Yu et al., 2024; Dong et al., 2024), but discrete sampling operations prevent gradient backpropagation. In contrast, CLaRa uniquely combines compression and joint training: by employing length-flexible compressed vectors in a shared latent space, we enable efficient differentiable selection while drastically reducing context length. Optimized solely through the generator’s language modeling loss, CLaRa ensures consistent training–inference alignment and efficient end-to-end learning without explicit retrieval supervision.
@@ 7 Conclusion
Impact Statement
This work aims to advance research in retrieval-augmented generation (RAG) by exploring compression-based representations that unify document understanding and generation. By enabling more compact and semantically dense representations, our approach has the potential to improve the efficiency and scalability of information access systems, particularly in settings where computational or memory resources are constrained.
From a societal perspective, improved retrieval and representation efficiency may benefit applications such as scientific search, education, and knowledge-intensive assistants by reducing latency and resource consumption while maintaining answer quality. At the same time, our current study relies on compressors pretrained primarily on Wikipedia data, which may reflect existing coverage biases in large-scale public corpora. Future work that incorporates more diverse and domain-specific data sources may help mitigate such limitations and improve robustness across modalities and application domains.
We do not identify any immediate negative ethical risks uniquely introduced by our method beyond those commonly associated with retrieval-based language models, such as potential biases inherited from training data or downstream misuse of generated content. As research progresses toward reasoning-oriented or agentic RAG systems that rely on compressed internal representations, it will be important to further examine transparency, controllability, and failure modes in complex decision-making settings.
Overall, this work contributes to ongoing efforts in machine learning to develop more efficient, generalizable, and reasoning-capable systems. We believe its broader societal impacts are aligned with well-established directions in the field and do not raise new ethical concerns that require special consideration at this stage.
@@
$\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y\,|\,x)}\Big[\underbrace{p(d\,|\,x)\big(p(y\,|\,x,d)-p(y\,|\,x)\big)}_{\text{(I) probability path}}+\underbrace{p(d\,|\,x)\,\frac{\partial p(y\,|\,x,d)}{\partial s_{xd}}}_{\text{(II) representation/generation path}}\Big].$ (A.7)
@@
$\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y\,|\,x)}\;p(d\,|\,x)\,\big(p(y\,|\,x,d)-p(y\,|\,x)\big).$ (A.9)
@@
$\frac{\partial p(y\,|\,x,d)}{\partial s_{xd}}=p(y\,|\,x,r)\;g^{\top}\frac{\partial r}{\partial s_{xd}}=\frac{p(y\,|\,x,r)}{\tau}\;\pi_{d}\;g^{\top}\big(z_{d}-r\big).$ (A.13)
@@
$\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y\,|\,x)}\Big[p(d\,|\,x)\,\big(p(y\,|\,x,d)-p(y\,|\,x)\big)\;+\;\frac{p(r\,|\,x)\,p(y\,|\,x,r)}{\tau}\;\pi_{d}\;g^{\top}\big(z_{d}-r\big)\Big].$ (A.14)